An aeroplane is flying at 1000 km/h through still air having a pressure of

78.5 kN/m2 (abs.) and temperature – 8°C. Calculate on the stagnation point on the nose of the

plane :

(i) Stagnation pressure,

(ii) Stagnation temperature, and

(iii) Stagnation density.

Take for air : R = 287 J/kg K and γ = 1.4


1
Expert's answer
2021-09-20T06:12:56-0400

The sonic velocity for adiabatic flow is given by "C_0=\sqrt{\gamma RT}" and then the Mach number as "M_0=\cfrac{V_0}{C_0}". These values need to be calculated because the stagnation pressure is given by


"p_s=p_0 {\Big[ 1+\Big( \dfrac{\gamma-1}{2} \Big){M_0}^2 \Big]}^{\dfrac{\gamma}{\gamma-1}}"


If we substitute the prior equation we have the following relation between all the variables we have and the stagnation pressure:

"p_s=p_0 {\Big[ 1+\Big( \dfrac{\gamma-1}{2\gamma RT} \Big){V_0}^2 \Big]}^{\dfrac{\gamma}{\gamma-1}}"


We proceed to substitute and calculate with

"p_0=78.5\,{kN/m^2}
\\T_0=-8+273=265 \,K
\\R=287\,J/kgK
\\ \gamma=1.4"

"V_0=\cfrac{1000\,km}{h}\times\cfrac{1000\,m}{1\,km}\times\cfrac{1\,h}{3600\,s}
\\ \implies V_0=277.77\,{m/s}"


and we find:

"p_s=(78.5\,kN/m^2) {\bigg[ 1+\Big( \dfrac{1.4-1}{2(1.4)(287\,{J/kgK})(265\,K)} \Big){(277.77\,{m/s})}^2 \bigg]}^{\cfrac{1.4}{1.4-1}}
\\ \text{ }
\\ \implies p_s=(78.5\,kN/m^2) {\Big[ 1.14492552 \Big]}^{3.5}
\\ \text{ }
\\ p_s=126.064\,{kN/m^2}"


(b) The stagnation temperature is given by


"T_s=T_0 {\Big[ 1+( \frac{\gamma-1}{2} ){M_0}^2 \Big]}
\\ \implies T_s=T_0{\Big[ 1+\Big( \dfrac{\gamma-1}{2\gamma RT} \Big){V_0}^2 \Big]}"


After the same substitution we can find the temperature:


"T_s=(265\,K) {\Big[ 1+\Big( \dfrac{1.4-1}{2(1.4)(287\,{J/kgK})(265\,K)} \Big){(277.77\,{m/s})}^2 \Big]}
\\ \text { }
\\ T_s=(265\,K) {\Big[ 1.14492552 \Big]} 
\\ \text { }
\\ \implies T_s=303.4\,K =30.4°C"


(c) At last, the stagnation density (ρs) is given by


"\cfrac{p_s}{\rho_s}=RT_s \implies \rho_s = \cfrac{p_s}{RT_s }"


We substitute the prior values and we find


"\rho_s = \cfrac{126064\,{N/m^2}}{(287\,{J/kgK})(303.4\,K)}
\\ \text { }
\\ \implies \rho_s =1.448\,{kg/m^3}"


In conclusion, (i) the stagnation pressure is "p_s=126.064\,{kN/m^2}" , while the (ii) stagnation temperature is "T_s=303.4\,K =30.4°C", and (iii) the stagnation density is "\rho_s =1.448\,{kg/m^3}".

Reference:

  • Rajput, R. K. (2005). A textbook of engineering thermodynamics. Laxmi Publications.